To identify a diatomic gas (X2), a researcher carried out the following experiment: She weighed an empty 3.9-L bulb, then filled it with the gas at 1.70 atm and 30.0  ∘C and weighed it again. The difference in mass was 7.5 g . Identify the gas.
Answer #1

Given:
P = 1.7 atm
V = 3.9 L
T = 30.0 oC
= (30.0+273) K
= 303 K

find number of moles using:
P * V = n*R*T
1.7 atm * 3.9 L = n * 0.08206 atm.L/mol.K * 303 K
n = 0.2666 mol

mass(solute)= 7.5 g


use:
number of mol = mass / molar mass
0.2666 mol = (7.5 g)/molar mass
molar mass = 28.13 g/mol

This is molar mass of N2
Answer: N2